OpenStudy (anonymous):

A circle has its center at the origin, and (5, -12) is a point on the circle. How long is the radius of the circle?

OpenStudy (kainui):

 Does the pythagorean theorem mean anything to you? Might it be potentially of use here?

OpenStudy (anonymous):

Is that way you solve it? I having a hard time understanding the course..

OpenStudy (kainui):

Yeah, just see if you can draw the picture and see how a circle's radius is just the hypotenuse of a right triangle.

OpenStudy (anonymous):

so 13?

OpenStudy (kainui):

Yep, perfect. =

OpenStudy (owlcoffee):

Let's first begin by interpretating the problem, recognizing it and of course, recalling what we know in order to find a solution. Now we are given a circumference and a point belonging to it, so we can use ANALYTIC GEOMETRY to solve it. Now, it has center O, meaningthat the coordinates are O(0,0) and the blonging point P(5,-12). The circle is defined by all the point that have the same distance to a point we call center (O, in this case), the distance all points have to thatcenter is called "radius". Now, I'll make a little proof for you, just because I just had breakfast and I'll leave soon: Let P1(x1,y1) and P2(x2,y2) arbitrary in the plane:  Now by pythagorean theorem, we say: \[d(P1,P2)^{2}=(x _{2}-x _{1})^{2}+(y _{2}-y _{1})\] \[d(P1,P2)=\sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\] and that's what we'll call the "distance between two points in the plane" formula. applying that little formula to the problem, replacing P1 as O(0,0) and P2 as P(5,-12): \[d(O,P)=\sqrt{(5-0)^{2}+(-12-0)^{2}}\] \[d(O,P=r\] PD: "r" states for "radius" so we can say: \[r=\sqrt{(5)^{2}+(-12)^{2}}\] \[r=\sqrt{25+144}\] \[r=\sqrt{169}\] \[r=13\] and with that, we have found the radius we were asked for in the exercise.